# Safe MPC Alignment with Human Directional Feedback

### v2

@@ III-C1 Representation of Human Intended Safety
Consider a human-robot coexistence task, where a human is supervising the robot motion. The user has his/her intent of robot motion safety. In Safe MPC (1), we hypothesize user-intended safety can be represented by a subset of the trajectory space:
@@
Here, $\boldsymbol{\theta}_{H}$ is intent parameter from an implicit set $\boldsymbol{\theta}_{H}\in\Theta_{H}\subset\mathbb{R}^{r}$. $\Theta_{H}$ is implicit to human and unknown to the robot. For the robot policy $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$, if $\boldsymbol{\theta}\in\Theta_{H}$, the user considers the robot motion to be safe; otherwise if $\boldsymbol{\theta}\notin\Theta_{H}$, the robot motion by $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$ might not be safe at certain time in the view of the user, and thus the user provides corrections to improve its safety. We use an implicit set $\Theta_{H}\in\mathbb{R}^{r}$, instead of single point $\boldsymbol{\theta}_{H}$, to capture the potential variation of user intents. We suppose that a volume measure of $\Theta_{H}$, $\operatorname{{Vol}}(\Theta_{H})$, is non-zero, and a user intent $\boldsymbol{\theta}_{H}\in\Theta_{H}$ can vary at different steps and context, but always within $\Theta_{H}$.
@@ III-C2 Human Directional Correction
We consider human gives correction to a robot in an online fashion while the robot is executing its most updated $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$. Specifically, let $i$ be the index of learning update, and suppose robot current policy is $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$. During rolling out $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the robot receives a sparse human directional correction $\boldsymbol{a}_{k_{i}}\in\mathbb{R}^{m}$ at the rollout time step $k_{i}$.
Here, we do not assume precise timing $k_{i}$ or duration of the human correction due to the sparsity or intermittency of human feedback. In other words, human can apply one or multiple corrections at any time in robot rollout time steps for any duration as he/she finds necessary. But we here assume human directional corrections $\boldsymbol{a}_{k_{i}}$ are in robot action space. For example, in autonomous driving where the robot action space is steering angle and acceleration, human can directly apply feedback with steering wheel and pedals. In some other applications, the action space correction can be achieved via specific mapping or interface. For example, to correct a torque-controlled robot arm, human’s correction on robot end-effector can be mapped into joint torque via Jacobian transpose matrices. The reason why we only consider the action space correction is because first, corrections in input spaces may be easier to implement, as shown in our later experiments, and second, corrections in state spaces can be infeasible for the under-actuated robot system [20].
After receiving each $\boldsymbol{a}_{k_{i}}$, the robot updates the safety parameter $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{i+1}$ in (1) for its MPC policy. Then, the robot continues with $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i+1}}^{\gamma}$ until the next possible user corrections. This process repeats until the convergence of the policy update or the user is satisfied with the robot motion. An illustration of one update is shown in Fig. 2.
@@
In the above learning process, we consider that a human user only provides directional correction $\boldsymbol{a}$, i.e., the magnitude $\left\lVert\boldsymbol{a}\right\rVert$ does not matter (theoretical justification will be given later). For example, when correcting a mobile robot, the user’s correction can be "left" or "right", without specifying how much the robot should go. As shown in our previous work [20] or later, the type of directional feedback provides a consistently large and flexible correction space, effectively overcoming the over-shooting correction, compared to existing magnitude feedback.
@@ III-C3 Robot Operation Safety and Emergency Stop
In the aforementioned learning procedure, the robot will respond to human corrections by executing the updated safe MPC policy $\pi_{\boldsymbol{\theta}}^{\gamma}$. However, it is possible that the robot may still exit the user-intended safety region. In such cases, a human user may trigger an emergency stop mechanism to stop the robot motion and potentially restart it from a desired safety region. In this sense, all trajectories before the human triggers the emergency stop are considered safe in human intent. In other words, until an emergency stop is triggered, human corrections typically occur while the robot is still within the intended safety zone; otherwise, the corrections would be useless. For instance, correcting a vehicle after it is crashed is futile.
@@ III-C4 Parametric Safety Constraints
In the safe MPC (4), the safety set is defined as a zero sublevel set of a safety value function ${g}_{\boldsymbol{\theta}}$. We note that if ${g}_{\boldsymbol{\theta}}$ has sufficient representation power, the safety set can represent complex (non-convex) safety region, which can be alternatively represented by the intersection of multiple constraints. In the following, we mainly focus on a weighted-features parameterization of ${g}_{\theta}$, which has been commonly used by the prior work on learning cost/reward [4, 20, 50, 51, 52]. Specifically, we parameterize $\boldsymbol{g}_{\theta}$
@@
with $\boldsymbol{\theta}=[\theta_{1},\theta_{2},...\theta_{r}]^{\mkern-1.5mu\mathsf{T}}\in\mathbb{R}^{r}$ and $\boldsymbol{\phi}=[\phi_{1},\phi_{2},\dots,\phi_{r}]^{\mkern-1.5mu\mathsf{T}}$ the weight and feature vectors, respectively. We fix the weight of $\phi_{0}$ as unit, because otherwise, $\boldsymbol{\theta}$ has scaling ambiguity, i.e., any scaled $c\cdot\boldsymbol{\theta}$ with $c>0$ lead to the same MPC solutions. The dimension (number) $r$ of features can be chosen such that $\boldsymbol{\phi}$ has enough expressiveness to represent the human-intended safety constraint. We do not impose any constraints on the choice of feature vector as long as it is differentiable, for example, $\boldsymbol{\phi}$ can be neural features. In practice the feature vector can be chosen by trial and error, i.e., starting from a few features and gradually adding more until the learned constraint is satisfactory to the human user. Later, we will discuss our proposed method for generic parameterization $\boldsymbol{g}_{\boldsymbol{\theta}}$.
@@ III-C5 Problem of Safe MPC Alignment
Following the procedure of human correction and robot policy update in Subsection III-C and given a weighted-feature safety constraint (5), we aim to develop a policy update rule: $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{i+1}$ for each received human directional correction, such that $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{H}\in\Theta_{H}$, as $i=1,2,3,\dots,$. i.e., safety constraints in the robot’s safe MPC policy will be aligned with human intent.
IV Overview of Safe MPC Alignment Algorithm
@@ IV-A Hypothesis on Human Directional Corrections
Informally, we have the following intuitions on human feedback. First, human corrections, on average, improve the robot safety. Second, human correction on robot safety is typically proactive and happens before the robot leaves the human-intended safety set. In other words, human correction is made when the robot is still in $\mathcal{S}(\boldsymbol{\theta}_{H})$ in (4). For example, it is useless to correct a vehicle after its crash. Third, human feedback is typically safety-margin aware: as a robot approaches the safety boundary, human users tend to prioritize safety over the task objective. For instance, when a vehicle moves closer to obstacles, humans may instinctively shift their focus to avoiding collisions rather than pursuing the original target. In the following, we will formalize those intuitions.
Without loss of generality, suppose the motion plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ is a local solution to the penalty MPC (3), and when the robot execute the first action $\boldsymbol{u}_{0}^{\boldsymbol{\theta},\gamma}$ (receding horizon implementation), a human applies a directional correction $\boldsymbol{a}$ in robot action space. Define
@@
We assume there is a nonzero volume set $\bar{\Theta}_{H}$ (containing a ball) in user implicit set $\Theta_{H}$, i.e., $\bar{\Theta}_{H}\subseteq\Theta_{H}$, such that for any $\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H}$, the human correction $\boldsymbol{a}$, on average, indicates a direction of lowering the penalty value $B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})$ for the robot motion plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$. Concretely, it means $\mathbf{a}$ on averages points in the direction of the negative gradient of $B(\boldsymbol{\xi},\boldsymbol{\theta}_{H})$ at $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, i.e.,
@@
[…]
(7) indicates that the magnitude of $\mathbf{a}$ does not affect the inequality. This brings one advantage of directional correction compared to the magnitude-based correction: regardless of robot plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, the allowable human corrections satisfying (7) always account for half of the action space, making directional corrections more accessible for human users, as shown in Fig. 3. Recall that $B(\boldsymbol{\xi},\boldsymbol{\theta})=J(\boldsymbol{\xi})-\gamma\ln(-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}))$, by extension of the gradient, (7) becomes
@@
Figure 3: Illustration of directional corrections in the space of robot motion plan $\boldsymbol{\xi}$s. The grey circles represent the level contours of $B(\boldsymbol{\xi},\boldsymbol{\theta}_{H})$. The black dot stands for current motion plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ and the red point stands for the plan $\boldsymbol{\xi}_{\boldsymbol{\theta}_{H}}^{\gamma}$ with human intended safety. The blue region stands for the half-space of all possible human corrections under our assumption. Some samples of human corrections are shown by blue arrows.
@@
(I) The first and second terms on the left side of (8) correspond to human correction toward lowering the cost and improving the safety of $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, respectively. $\gamma$ serves as a weight between two improvements.
(II) If we interpret the value of $-{g_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})}$ as the safety margin of $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, (8) suggests that the human correction is safety-margin aware: the closer the robot motion get to the unsafe region, the more will human user emphasize on improving the robot’s safety than minimizing the objective. This property is consistent with our previous intuition.
@@
This corresponds to the third intuition: a user typically gives correction only when the robot is in user intended safety region, i.e., $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}\in\mathcal{S}(\boldsymbol{\theta}_{H})$.
In sum, we have used (7) and (9) to formalize the intuitions we have observed at the beginning of this session. Those two constraints on $\boldsymbol{\theta}_{H}$ will be used in the robot learning next.
IV-B Alignment Algorithm: Learning by Hypothesis Space Cutting
To find $\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H}\subseteq\Theta$ for the safe MPC (1), our idea is that for each human directional correction received $i=0,1,2,....$, the robot performs one learning iteration, by updating a hypothesis space $\Theta_{i}\in\mathbb{R}^{r}$, a set containing all possible $\bar{\Theta}_{H}$. At initial, the robot may choose a large enough hypothesis space $\Theta_{0}$. Specifically, one can choose $\Theta_{0}$ as a box space
@@ IV-B Alignment Algorithm: Learning by Hypothesis Space Cutting
in which $\boldsymbol{c}_{l},\boldsymbol{c}_{u}$ are the vectors of lower and upper bounds of the box space, resepctively, and the inequalities applied element-wise. By saying a $\Theta_{0}$ large enough, we assume
@@
Later, we will show that our method can self-access the learning failure in case of misspecification, where (11) is not met.
Recall the procedure of human correction in Section III-C and Fig. 2. At $i$-th learning iteration, $i=1,2,3...$, the robot executes $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ with $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$. The user gives a directional correction $\boldsymbol{a}_{k_{i}}$ at rollout step ${k_{i}}$. By letting $\mathbf{\bar{a}}_{i}=(\boldsymbol{a}_{k_{i}}^{T},0,0,...)$ as in (6), we can establish the following two constraints on $\bar{\Theta}_{H}$ from the previous hypotheses (7) and (9):
$\bar{\Theta}_{H}\subset\mathcal{C}_{i}=\left\{\boldsymbol{\theta}|\left<\mathbf{\bar{a}}_{i},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta})\right>\leq 0,\,\,\,g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<0\right\},$ (12)
which we refer to as the cutting set. Note that the first inequality $\mathbf{\bar{a}}_{i},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta})\right>\leq 0$ is deduced by multiplying -1 on both sides of (7). Using $\mathcal{C}_{i}$, one can update the hypothesis space
@@
The above update can be viewed as a step of using (12) to cut (intersect with) the current hypothesis space $\Theta_{i-1}$ to remove the portion where $\bar{\Theta}_{H}$ is impossible. The remaining hypothesis space $\Theta_{i}$ will be new hypothesis used for the next iteration. With the above update, we have the following result.
@@
As shown in Fig. 4, the procedure of robot learning is to cut and move hypothesis space based on human directional correction. Therefore, we name our alignment method Hypothesis Space Cutting. As a summary, we present the three main steps of our algorithm below.
Figure 4: The update of constraint parameter (2D examples). Two cutting lines are used to cut $\Theta_{i-1}$ to $\Theta_{i}$. The user implicit intent set $\bar{\Theta}_{H}$ is contained in $\Theta_{i}$. New $\boldsymbol{\theta}_{i+1}$ is chosen as the parameter for the next safe MPC policy.
Safe MPC alignment by Hypothesis Space Cutting Initialize $\Theta_{0}$ that contains $\bar{\Theta}_{H}$. Repeat the following three steps for $i=1,2,...,K$ until $\boldsymbol{\theta}_{i}\in\bar{\Theta}_{H}$. Step 1: Robot chooses parameter $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ for its current safe MPC policy $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$. Step 2: Human user applies a directional correction $\boldsymbol{a}_{k_{i}}$ while robot executes $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ Step 3: Perform hypothesis space cut using (13) to get new hypothesis space $\Theta_{i}$.
At $i$th learning iteration, the robot chooses $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ from the most recent hypothesis space $\Theta_{i-1}$ for its current MPC $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$. While executing $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the robot receives human directional correction $\boldsymbol{a}_{k_{i}}$ at the rollout time step $k_{i}$. This correction leads to a cutting set $\mathcal{C}_{i}$ (12), which will be used to update $\Theta_{i-1}$ to $\Theta_{i}$ in (13). The above iteration repeats $i=1,2,3,...$. From (13), one can easily see that $\Theta_{0}\supseteq\Theta_{1}\supseteq...\supseteq\bar{\Theta}_{H}$. Given a volume measure, it follows $\operatorname{{Vol}}\Theta_{0}\geq\operatorname{{Vol}}\Theta_{1}\geq...,\geq\operatorname{{Vol}}{\bar{\Theta}_{H}}$. Thus, the volume of the hypothesis space is at least not increasing after each user correction. It is hoped that the volume of the hypothesis space decreases quickly, such that new $\boldsymbol{\theta}_{i+1}$ increases its chance of falling inside $\bar{\Theta}_{H}$ — a successful safety alignment between the robot and human.
For the above Safe MPC Alignment method, we will later (in Section V) show that within finite number $K$ of corrections, the method will either

### v3

@@ III-C1 Representation of Human Intended Safety
Consider a human is supervising the robot’s behavior. In Safe MPC (1), we hypothesize that user-intended safety can be represented by a subset of the trajectory space:
@@
Here, $\boldsymbol{\theta}_{H}$ is from a set $\boldsymbol{\theta}_{H}\in\Theta_{H}$, encoding all possible user intended safety. $\Theta_{H}$ is implicit and unknown to the robot. For the robot policy $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$, if $\boldsymbol{\theta}\in\Theta_{H}$, the user considers the planned robot motion to be safe; otherwise if $\boldsymbol{\theta}\notin\Theta_{H}$, the robot motion by $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$ might not be safe at certain time. We use an implicit set $\Theta_{H}\in\mathbb{R}^{r}$, instead of a single point $\boldsymbol{\theta}_{H}$, to capture the potential variance of user intents. We suppose that a volume measure of $\Theta_{H}$, $\operatorname{{Vol}}(\Theta_{H})$, is non-zero, and a user intent $\boldsymbol{\theta}_{H}\in\Theta_{H}$ can vary at different steps and contexts, but always within $\Theta_{H}$.
@@ III-C2 Human Directional Correction
As shown in Fig. 2, we consider human correcting a robot online while the robot is executing $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$. Specifically, let $i$ be the index of learning update, and suppose robot current policy is $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$. During rollout of $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the robot receives a human correction $\boldsymbol{a}_{k_{i}}\in\mathbb{R}^{m}$ at the rollout time step $k_{i}$. Here, we do not pose any assumption on correction time $k_{i}$ and its duration due to the sparsity or intermittency of human feedback. In other words, the human can apply one or multiple corrections at any time as necessary. But we assume human correction $\boldsymbol{a}_{k_{i}}$ is in the robot action space. For example, in autonomous driving where the robot action space is steering angle and acceleration, human can directly apply feedback with the steering wheel and pedals. In some other applications, the action space correction can be achieved via a specific user interface. For example, to correct a torque-controlled robot arm, human’s correction on the end-effector can be mapped into joint torque via the Jacobian transpose. The reason why we only consider the action space correction is because first, corrections in input spaces may be easier to implement, as shown in our later experiments, and second, corrections in state spaces can be infeasible for the under-actuated robot system.
After receiving a human correction $\boldsymbol{a}_{k_{i}}$, the robot updates the safety parameter $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{i+1}$ in (1) for its Safe MPC. Then, the robot continues with $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i+1}}^{\gamma}$ until the next possible user correction. This process repeats until the convergence of the policy update or the user is satisfied with the robot’s motion.
@@
While the correction $\boldsymbol{a}$ given by human is real-valued, only its direction is used in the learning process. This means that we will use the sign (for scalar signal) or its normalized unit vector (for vector-valued correction signal) for learning update. The magnitude of the correction is NOT taken into account in the learning process. As we will show later, directional feedback provides a consistently large and flexible correction space, effectively overcoming the over-shooting correction, compared to magnitude feedback.
@@ III-C3 Robot Operation Safety and Emergency Stop
In the aforementioned learning procedure, the robot will respond to human correction by executing the updated safe MPC policy $\pi_{\boldsymbol{\theta}}^{\gamma}$. However, it is possible that the robot may still exit the user-intended safety. In such cases, the human may trigger an emergency stop to stop the robot and potentially restart it from a desired safety region. In such case, the executed motion before the emergency stop is assumed safe. In other words, until an emergency stop is triggered, human correction typically occur when the robot is still in the intended safety set; otherwise, the correction would be useless. For instance, correcting a vehicle after it is crashed is futile.
@@ III-C4 Parametric Safety Constraints
In (4), the safety set is defined as a zero sublevel set of ${g}_{\boldsymbol{\theta}}$. If ${g}_{\boldsymbol{\theta}}$ has sufficient representation power, the safety set can represent a complex safety region. In the following, we focus on a weighted-features parameterization of ${g}_{\boldsymbol{\theta}}$, which has been commonly used by the prior work on learning cost/reward [lindner2023learning, jin2022learning, zhang2023dual, biyik2022aprel, bobu2020quantifying]. Specifically,
@@
with $\boldsymbol{\theta}=[\theta_{1},\theta_{2},...\theta_{r}]^{\mkern-1.5mu\mathsf{T}}\in\mathbb{R}^{r}$ and $\boldsymbol{\phi}=[\phi_{1},\phi_{2},\dots,\phi_{r}]^{\mkern-1.5mu\mathsf{T}}$ the learnable weight and feature vectors, respectively. We fix the weight of $\phi_{0}$ as unit, because otherwise, $\boldsymbol{\theta}$ has scaling ambiguity, i.e., any scaled $c\cdot\boldsymbol{\theta}$ with $c>0$ lead to the same constraint. In this work, we assume that the features in $\boldsymbol{\phi}$ are given as a priori. These features, defined in system states (e.g., position, velocity) and actions, can be either manually defined or obtained from pre-trained neural representations, leveraging prior knowledge of the environment and tasks, as demonstrated in our experiments. The feature selection itself is beyond the scope of this paper and is left for future work. A more detailed discussion of this assumption and its limitations is in Section IX-A3.
@@ III-C5 Problem of Safe MPC Alignment
Following the procedure of human correction and robot policy update in Section III-C, we aim to develop a policy update rule: $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{i+1}$ for each received human correction, such that $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{H}\in\Theta_{H}$, as $i=1,2,3,\dots,$. i.e., safety constraints in the robot’s safe MPC policy will be aligned with human intent.
IV Overview of Safe MPC Alignment
@@ IV-A Hypothesis on Human Directional Corrections
Informally, we have the following intuitions on human feedback. First, human corrections, on average, improve the robot’s safety. Second, human correction is typically proactive and happens before the robot leaves the human-intended safety set. In other words, human correction is made when the robot is still in $\mathcal{S}(\boldsymbol{\theta}_{H})$ in (4). For example, it is useless to correct a vehicle after its crash. Third, human feedback is typically safety-margin aware: as a robot approaches the safety boundary, human users tend to prioritize safety over the task cost. For instance, when a vehicle moves closer to obstacles, humans may instinctively shift their focus to avoiding collisions rather than pursuing the original target. In the following, we will formalize those intuitions.
Without loss of generality, suppose the motion plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ is a local solution to the penalty MPC (3), and when the robot execute the first action $\boldsymbol{u}_{0}^{\boldsymbol{\theta},\gamma}$ (receding horizon implementation), a human applies a directional correction $\boldsymbol{a}$ in robot action space. In this subsection, the subscription $k_{i}$ is omitted for notation simplicity. Define
@@
We assume there is a nonzero volume set $\bar{\Theta}_{H}$ (containing a ball) in user implicit set $\Theta_{H}$, i.e., $\bar{\Theta}_{H}\subseteq\Theta_{H}$, such that for any $\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H}$, the human correction $\mathbf{a}$, on average, indicates a direction of lowering the penalty value $B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})$. Concretely, it means $\mathbf{a}$ on averages points in the direction of the negative gradient of $B(\boldsymbol{\xi},\boldsymbol{\theta}_{H})$ at $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, i.e.,
@@
Here, $\left\langle\cdot,\cdot\right\rangle$ denotes the inner product operation, and ${\operatorname{\mathbb{E}}}_{\mathbf{a}}(\mathbf{a})=\mathbf{\bar{a}}$ denotes the expectation value of $\mathbf{a}$. $\nabla B(\boldsymbol{\xi},\boldsymbol{\theta})$ denotes the full gradient of $B$ with respect to $\boldsymbol{u}_{0:T-1}$ evaluated at trajectory $\boldsymbol{\xi}=\{\boldsymbol{u}_{0:T-1},\boldsymbol{x}_{0:T}\}$. Similar notations apply to $J$ and $g$ below. Since $\nabla B(\boldsymbol{\xi},\boldsymbol{\theta})$ is not dependent on $\mathbf{a}$, the expectation can be moved inside the inner product. Because state sequence $\boldsymbol{x}_{0:T}$ is the rollout of $\boldsymbol{u}_{0:T-1}$ on dynamics, the differentiation goes through the dynamics equation.
Equation (7) shows that although the human correction $\boldsymbol{a}$ is provided as a real-valued vector, only its direction—not its magnitude—affects the inequality. This brings one advantage of directional correction compared to the magnitude-based correction: the allowable human correction satisfying (7) always account for half of the action space, making directional corrections more accessible for human users, as shown in Fig. 3. Recall that $B(\boldsymbol{\xi},\boldsymbol{\theta})=J(\boldsymbol{\xi})-\gamma\ln(-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}))$, by extension of the gradient, (7) becomes
@@
Figure 3: Illustration of directional correction in the space of robot plan $\boldsymbol{\xi}$s. The grey circles represent the level contours of $B(\boldsymbol{\xi},\boldsymbol{\theta}_{H})$. The black dot stands for the current motion plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ and the red point stands for optimal plan $\boldsymbol{\xi}_{\boldsymbol{\theta}_{H}}^{\gamma}$ with $\theta_{H}$. The blue region stands for the half-space of all possible human correction under our assumption (7). Some samples of human corrections are shown by blue arrows.
@@
(I) The first and second terms of (8) correspond to human correction toward lowering the cost and improving the safety of $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, respectively. $\gamma$ weighs between two improvements.
(II) If we interpret the value of $g_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})$} as the safety margin of the plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, (8) suggests that the human correction is safety-margin aware: the closer the robot motion get to the unsafe region, the more will human improve the robot safety than minimizing the cost. This property is consistent with our previous intuition.
@@
This corresponds to the third intuition: a user typically gives correction only when the robot is in user intended safety region.
In sum, the two constraints(7) and (9) formalize the intuitions we have observed at the beginning of this subsection. Those constraints will be used in the robot learning next.
IV-B Main Algorithm: Learning by Hypothesis Space Cutting
To find $\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H}\subseteq\Theta$ for the Safe MPC, our idea is that for each human correction $i=0,1,2,.$..., the robot performs one learning iteration, by updating a hypothesis space $\Theta_{i}\in\mathbb{R}^{r}$, a set containing all possible $\bar{\Theta}_{H}$. Initially, the robot may choose a large box hypothesis space $\Theta_{0}$:
@@ IV-B Main Algorithm: Learning by Hypothesis Space Cutting
in which $\boldsymbol{c}_{l},\boldsymbol{c}_{u}$ are the vectors of lower and upper bounds of the box space, respectively, and the inequalities are applied element-wise. By saying a $\Theta_{0}$ large enough, we assume
@@
Later, we will show our method can self-access the misspecification, where (11) is not met.
[…]
$\mathcal{C}_{i}=\left\{\boldsymbol{\theta}|\left<\mathbf{\bar{a}}_{i},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta})\right>\leq 0,\,\,\,g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<0\right\},$ (12)
which we refer to as the cutting set. Using $\mathcal{C}_{i}$, one can update the hypothesis space
@@
The above update can be viewed as using (12) to cut the hypothesis space $\Theta_{i-1}$ to remove portions where $\bar{\Theta}_{H}$ is impossible. The new hypothesis space $\Theta_{i}$ will be used for the next iteration. With the above update, we have the following result.
@@
As in Fig. 4, the procedure of robot learning is to cut and move the hypothesis space based on human directional correction. Therefore, we name our learning method Safe MPC Alignment by Hypothesis Space Cutting. As a summary, we present the three main steps of our algorithm below.
Figure 4: The update of hypothesis space (2D example). Two cutting lines are used to cut $\Theta_{i-1}$ to $\Theta_{i}$. The user intent set $\bar{\Theta}_{H}$ is contained in $\Theta_{i}$. New $\boldsymbol{\theta}_{i+1}$ is chosen from $\Theta_{i}$ for the updated Safe MPC policy.
Safe MPC Alignment by Hypothesis Space Cutting Initialize $\Theta_{0}$ that contains $\bar{\Theta}_{H}$. Repeat the following three steps for i=1,2,,Ki=1,2,...,K until $\boldsymbol{\theta}_{i}\in\bar{\Theta}_{H}$. Step 1: Robot chooses parameter $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ for its current Safe MPC $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$. Step 2: Human user applies a directional correction $\boldsymbol{a}_{k_{i}}$ while robot executes $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ Step 3: Perform hypothesis space cut using (13) to get new hypothesis space $\Theta_{i}$.
At $i$th learning update, the robot chooses $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ for its current Safe MPC $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$. While executing $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the robot receives human correction $\boldsymbol{a}_{k_{i}}$ at the rollout time step $k_{i}$. This correction leads to a cutting set $\mathcal{C}_{i}$ (12), which will update $\Theta_{i-1}$ to $\Theta_{i}$ in (13). The above iteration repeats $i=1,2,3,$.... From (13), one can easily see that $\Theta_{0}\supseteq\Theta_{1}\supseteq...\supseteq\bar{\Theta}_{H}$. Given a volume measure, it follows $\operatorname{{Vol}}\Theta_{0}\geq\operatorname{{Vol}}\Theta_{1}\geq...,\geq\operatorname{{Vol}}{\bar{\Theta}_{H}}$. Thus, the volume of the hypothesis space is at least not increasing after each correction. It is hoped that the volume of the hypothesis space decreases quickly, such that new $\boldsymbol{\theta}_{i+1}$ increases its chance of choosing from $\bar{\Theta}_{H}$ — meaning a successful safety alignment between the robot and human.
For the Safe MPC Alignment, we will later (Section V) show within finite number $K$ of correction, the method will either